OpenStudy (anonymous):

The hypotenuse of a right triangle is 41 in long. One leg of the same triangle is 40 in long. What is the length of the other leg? Explain how you found your answer.

OpenStudy (anonymous):

PLEASE HELP ? I know it's 40 but how do I explain it ??

OpenStudy (zzr0ck3r):

a^2+b^2 = c^2 c = hypotenuse, a = leg, b = other leg c = 41 a = 40 so 40^2+b^2 = 41^2 b^2 = 41^2-40^2 = 81 so b^2 = 81 so b = sqrt(81) we only take the positive root because negative would not make sense.

OpenStudy (nincompoop):